Solve the following problems. Use the properties of the sampling distribution of

sample means.


5. If a population with size 3 has a standard deviation of 2.4, what is thestandard 

deviation of the sampling distribution of its means? The sampling distribution 

has a sample size of 2 and all possible samples are drawn without

replacements.



1
Expert's answer
2021-05-07T11:09:01-0400

Solution:

given, population size: N=3

"\sigma" = 2.4

sample size: n=2

Hence,

the standard deviation of the sampling distribution of its means= SD("\bar x" )="\dfrac{\sigma}{\sqrt n}" ="\dfrac{2.4}{\sqrt 2}" =1.6971


The standard deviation of the sampling distribution of sample mean=1.6971